OpenStudy (anonymous):

the probability of guessing the first three multiple choice exam each of which has four possible answers

OpenStudy (tkhunny):

What is p(guessing correctly on the first one)?

OpenStudy (phi):

on the first question, if there are 4 choices and you randomly pick one of them, what is the chance you got it correct? any idea ?

OpenStudy (dmndlife24):

1/4 probability you got one correct... Since it is guessed three times you would multiply 1/4 by itself 3 times \[\frac{ 1 }{ 4 } \times \frac{ 1 }{ 4 } \times \frac{ 1 }{ 4 } = ?\]

OpenStudy (tkhunny):

Here we are, again. Now, we'll never know if @robv23 can do it.

OpenStudy (dmndlife24):

?